Circular No. 7052

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1998 W1 T. B. Spahr, University of Arizona, reports his discovery of another

comet with the 0.41-m f/3 Schmidt telescope in the course of the Catalina Sky Survey. There is no visible tail, and the round coma is 18′′ in diameter.

1998 UT  α_{2000}  δ_{2000}  m_{1}
Nov. 16.38714  4^{h}33^{m}23.^{s}01  −10^{o}36′32′′.7
16.39505  4 33 22.71  −10 36 29.4
16.40318  4 33 22.31  −10 36 25.8  16.5
17.31281  4 32 43.95  −10 29 41.1
17.31659  4 32 43.79  −10 29 39.2

LEONID METEORS 1998 Early reports indicate that there was substantial Leonid activity earlier

than expected, with a high proportion of very bright fireballs, some brighter than mag −8, with enduring trains. Visual Leonid rates increased from ∼ 10/hr around Nov. 16.4 UT (G. Zay, Descanso, CA) to ∼ 140–150/hr around Nov. 16.9–17.1 (M. Langbroeck and H. Zhao, Delingha, Qinghai, China; and K. Hornoch, Lelekovice, Czech Republic). A. Pereira (Cabo da Roca, Portugal) observed 214 Leonids during Nov. 17.13–17.24. However, A. Fitzsimmons reports that a group of observers at La Palma observed ∼ 1000 Leonids/hr around Nov. 17.15, increasing to ∼ 2000/hr around Nov. 17.19. J. Gluckman, D. Healy, and others near Sierra Vista, AZ, reported 423 Leonids during Nov. 17.36–17.45, while A. Hale (Las Cruces, NM) reported only 126 Leonids during Nov. 17.38–17.47.

COMET C/1998 M5 (LINEAR) Visual m_{1} estimates: Oct. 20.83 UT, 10.5 (A. Kammerer, Ettlingen,

Germany, 0.20-m reflector); Nov. 7.76, 10.4 (R. Haver, Frasso Sabino, Rieti, Italy, 0.15-m reflector); 12.99, 10.4 (J. Bortle, Stormville, NY, 0.41-m reflector); 16.85, 10.9 (C. Vitorino, Cabo da Roca, Portugal, 0.25-m refl.).

COMET C/1995 O1 (HALE-BOPP) Visual m_{1} estimates: Oct. 13.53 UT, 9.9 (D. A. J. Seargent, Cowra,

N.S.W., 10×50 binoculars); 20.80, 10.5 (A. Pearce, Nedlands, W. Australia, 0.41-m reflector); Nov. 15.07, 10.8, 1’.5 (J. G. S. Aguiar, Joaquim Egidio, Brazil, 0.23-m reflector).

1998 November 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green